# Correlated electronic structures and unconventional superconductivity in bilayer nickelate heterostructures

### v1

@@ Abstract
The recent discovery of ambient-pressure superconductivity in thin-film bilayer nickelates opens new possibilities for investigating electronic structures in this new class of high-transition temperature ($\bm{T_{\mathrm{C}}}$) superconductors. Here, we construct a realistic multi-orbital Hubbard model for the thin-film system based on structural parameters integrating scanning transmission electron microscopy (STEM) measurements and ab initio calculations. The interaction parameters are calculated with the constrained random phase approximation (cRPA). The on-site Coulomb repulsive $\bm{U}$ and electron filling $\bm{n}$ dependencies of the correlated electronic structure are investigated based on density functional theory (DFT) plus cluster dynamical mean-field theory (CDMFT) methods. The fermi surface topology is substantially changed by $\bm{U}$ around the cRPA value, suggesting the indispensable role of correlation effects. A modified random-phase-approximation (RPA) approach, using $\bm{U=3.5}$ and $\bm{n=1.3}$, yields a pronounced $\bm{s^{\pm}}$-wave pairing instability, due to the strong spin fluctuations originated from Fermi surface nesting between bands with predominantly $\bm{d_{z^{2}}}$ characters. Our findings may explain the similar onset $\bm{T_{\mathrm{C}}}$ comparing pressured bulks and strained thin films of the bilayer nickelates.
@@
##\##These two authors contribute equally.
@@ Introduction
The high-temperature superconductivity in the Ruddlesden-Popper (RP) nickelates has garnered significant attention[1]. Extensive experimental[2, 3, 4, 5, 6, 7, 8, 9, 10] and theoretical[11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32] studies have been conducted to investigate their physical properties and superconducting mechanisms. The high-pressure conditions required for the superconducting phase in bulk compounds pose significant challenges for detailed property measurements[9]. Recently, the discovery of ambient-pressure superconductivity in bilayer nickelate epitaxial thin films grown on SrLaAlO4 substrates has opened new possibilities for experimental investigation[33, 34]. Structurally, the out-of-plane (i.e. c-axis) lattice constants in compressively strained films are elongated, opposite to anisotropically pressured bulks, raising intriguing question about the role of the interlayer coupling via the dz2 orbital in superconductivity.
On the theoretical front, understanding the electronic structure is essential for exploring the superconducting mechanism of these materials. In this work, we study the correlated electronic structure based on structural parameters from scanning transmission electron microscopy (STEM) experiments, in combination with first-principles calculations and many-body cluster dynamical mean-field theory (CDMFT) methods. To consider the effect of on-site Coulomb repulsion and the hole doping originated from Sr substitution of La observed in thin films[34], we study the $U$ and doping dependence of the electronic structure with DFT+CDMFT approach. Our results demonstrate a pronounced change of Fermi surfaces for an intermediate interaction strength near the cRPA parameter regime. A modified RPA method gives rise to a leading $s^{\pm}$-wave pairing symmetry, originated from the Fermi surface nesting between bands with $d_{z^{2}}$ characters.
Structural Analysis
@@
We begin with STEM analysis of oxygen octahedral rotations in a superconducting La2.85Pr0.15Ni2O7 film epitaxially grown on SrLaALO4 substrate[34]. In ambient-pressure bulk La3Ni2O7 crystals, the $a^{-}a^{-}c^{0}$ oxygen octahedral rotation pattern is generally seen (Fig. 1a): the NiO6 octahedra undergo inverse rotations along the $a$ and $b$ axes (i.e. [100]pc and [010]pc), while no rotation occurs along the $c$ axis ([001]pc). Note that [100]pc and [010]pc are symmetric directions under the symmetry group. These rotations lead to splittings of oxygen positions along the [100]pc/[010]pc, and [110]pc directions. In contrast, the La2.85Pr0.15Ni2O7 film grown on SrLaAlO4 appears more likely to conform to the high-symmetry $a^{0}a^{0}a^{0}$ pattern (Fig. 1b), for reasons to be revealed shortly. This pattern features the absence of splittings of the oxygen positions in any directions. Fig. 1c and d present zoomed-in annular bright-field (ABF) images of the La2.85Pr0.15Ni2O7 film cross-sections, projected along the [100]pc and [110]pc directions, respectively (larger scale images in Fig. S1). In the [100]pc direction, the oxygen atoms in both the experimental and simulated film results appear rounder compared to the elliptical shape observed in the simulated bulk. In the [110]pc direction, the oxygen atoms exhibit a distinct zigzag feature in the Ni-O layer for the bulk simulated case, which is not present in the film measurement and the film simulation. The line profiles of oxygen atoms (Fig. 1e) more clearly demonstrate their alignment in contrast to the bulk simulations. These results suggest that the coherent epitaxy may suppress the oxygen octahedral rotations in the film, giving rise to a higher symmetry of the lattice, thus further altering orbital overlap and bonding interactions between Ni and O atoms.
@@
The crystal structure of thin films is then constructed, using realistic structural parameters[34] (see Table 1), for DFT+$U$ calculations. A simple yet effective choice is to use the half unit-cell (UC) containing one bilayer of La3Ni2O7 (see Fig. 2a). Experiments show that near the interface, Sr substitution of La due to interfacial diffusion from the substrate[34] may introduce moderate hole doping into the system. Such a Sr-doping effect is simulated by the virtual crystal approximation, and the resultant band structure can be approximated by the rigid-band shift of the Fermi level in the one without Sr. More details of DFT+$U$ calculations can be found in supplementary materials (SM.2). The band structure is shown in Fig. 2b. Similar to the bulk structure[11, 35], the low-energy bands are mainly formed by the Ni-$e_{g}$ orbitals as demonstrated by the orbital-projected bands (red/blue solid circles for Ni-$d_{z^{2}}$/$d_{x^{2}-y^{2}}$ orbital).

### v2

@@ Abstract
The recent discovery of ambient-pressure superconductivity in thin-film bilayer nickelates opens new possibilities for investigating electronic structures in this new class of high-transition temperature ($\bm{T_{\mathrm{C}}}$) superconductors. Here, we construct a realistic multi-orbital Hubbard model for the thin-film system, by integrating ab initio calculations with scanning transmission electron microscopy (STEM) measurements, which reveal a higher-symmetry lattice. The interaction parameters are calculated with the constrained random phase approximation (cRPA). Density functional theory (DFT) plus cluster dynamical mean-field theory (CDMFT) calculations, with cRPA calculated on-site Coulomb repulsive $\bm{U}$ and experimentally measured electron filling $\bm{n}$, quantitatively reproduces Fermi surfaces from angle-resolved photoemission spectroscopy (ARPES) experiments. The distinct Fermi surface topology from simple DFT+$U$ results features the indispensable role of correlation effects. Based upon the correlated electronic structures, A modified random-phase-approximation (RPA) approach yields a pronounced $\bm{s^{\pm}}$-wave pairing instability, due to the strong spin fluctuations originated from Fermi surface nesting between bands with predominantly $\bm{d_{z^{2}}}$ characters. Our findings highlight the quantitative effectiveness of the DFT+cRPA+CDMFT approach that precisely determines correlated electronic structure parameters without fine-tuning. The revealed intermediate correlation effect may explain the same order-of-magnitude onset $\bm{T_{\mathrm{C}}}$ observed both in pressured bulk and strained thin film bilayer nickelates.
@@
School of Physics, Beijing Institute of Technology, Beijing 100081, China
@@
##\##These authors contribute equally.
@@ Introduction
The high-temperature superconductivity under pressure in the Ruddlesden-Popper (RP) nickelates has garnered significant attention[1]. Extensive experimental[2, 3, 4, 5, 6, 7, 8, 9, 10] and theoretical[11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32] studies have been conducted to investigate their physical properties and superconducting mechanisms. The high-pressure conditions required for the superconducting phase in bulk compounds pose significant challenges for detailed property measurements[9]. Recently, the discovery of ambient-pressure superconductivity in bilayer nickelate epitaxial thin films grown on SrLaAlO4 substrates open new possibilities for experimental investigations[33, 34, 35]. Structurally, the out-of-plane (i.e. $c$-axis) lattice constants in compressively strained films are elongated, opposite to anisotropically pressured bulks despite their similar onset $T_{\mathrm{C}}$, raising an intriguing question about the role of the interlayer coupling via the $d_{z^{2}}$ orbital in superconductivity.
On the theoretical front, understanding the electronic structure is essential for exploring the superconducting mechanism of these materials. There has been considerable debate on the electronic structure and the role played by the correlation effect in superconducting bilayer nickelate bulks under high pressure[16, 21, 36, 12, 5, 26, 30]. The superconducting bilayer nickelate thin films under ambient pressure enables angle-resolved photoemission spectroscopy (ARPES) experiments that serve as a benchmark to address this issue[37]. In this work, we study the correlated electronic structure and the pairing mechanism of bilayer nickelate superconducting thin films, based on structural parameters from experiments, and achieve quantitative agreements with ARPES Fermi surfaces without parameter fine-tuning.
Experimental Structural Analysis
@@
Scanning transmission electron microscopy (STEM) analysis of oxygen octahedral rotations confirms a higher-symmetry lattice in the superconducting bilayer nickelate epitaxial films grown on SrLaALO4 substrates[34], than that in the ambient-pressure bulks. In La3Ni2O7 bulk crystals without pressures, the $a^{-}a^{-}c^{0}$ oxygen octahedral rotation pattern is generally seen (Fig. 1a): the NiO6 octahedra undergo inverse rotations along the $a$ and $b$ axes (i.e. [100]pc and [010]pc), while no rotation occurs along the $c$ axis ([001]pc). Note that [100]pc and [010]pc are symmetric directions under the symmetry group. This rotation leads to splittings of oxygen positions along the [100]pc/[010]pc, and [110]pc projections. In contrast, the La2.85Pr0.15Ni2O7 film grown on SrLaAlO4 appears more likely to conform to the high-symmetry $a^{0}a^{0}a^{0}$ pattern (Fig. 1b). This pattern features the absence of splittings of the oxygen positions in any projection directions. Fig. 1c and d present zoomed-in annular bright-field (ABF) images of the La2.85Pr0.15Ni2O7 film cross-section, projected along the [100]pc and [110]pc directions, respectively (larger field of view images in Fig. S1). In the [100]pc direction, the oxygen atoms in both the experimental and simulated film results appear rounder compared to the elliptical shape observed in the simulated bulk. In the [110]pc direction, the oxygen atoms exhibit a distinct zigzag feature in the Ni-O layer for the bulk simulated case, which is not present in the film measurement and the film simulation. The line profiles of oxygen atoms (Fig. 1e) more clearly demonstrate their alignment in contrast to the bulk simulations. These results suggest that the coherent epitaxy may suppress the oxygen octahedral rotations in the film, giving rise to a higher symmetry of the lattice, thus further altering orbital overlap and bonding interactions between Ni and O atoms. Detailed positions of each cation in the lattice are measured combining STEM and X-ray diffraction (XRD) experiments, as shown in (Table 1).
@@
The crystal structure of thin films is constructed for DFT+$U$ calculations, using realistic structural parameters (Table 1). A simple yet effective choice is to use the half unit cell (UC) containing one nickelate bilayer (see Fig. 2a). Experiments show that Sr substitution of La near the interface due to interfacial diffusion from the substrate[34] may introduce moderate hole doping into the system. Such a Sr-doping effect is simulated by the virtual crystal approximation, and the resultant band structure can be approximated by the rigid-band shift of the Fermi level in the one without Sr. More details of DFT+$U$ calculations can be found in supplementary materials (SM.2). The band structure is shown in Fig. 2b. Similar to the bulk structure[11, 36], the low-energy bands are mainly formed by the Ni-$e_{g}$ orbitals as demonstrated by the orbital-projected bands.
A realistic multi-orbital Hubbard model for Ni-$e_{g}$ subspace is formulated (see Method). The first part is the tight-binding model which has bands aligned well with the DFT+$U$ band (see Fig. 2b). The TB model Hamiltonian in ${\bf k}$-space has the same matrix structure as in Ref. [11] The onsite energy level $\epsilon_{x/z}$ and the hopping parameters are summarized in Table 2. Here, we introduce the inter-layer bonding (+) and anti-bonding (-) orbital basis $d_{z_{\pm}}=(d_{Az}\pm d_{Bz})/\sqrt{2}$ and $d_{x_{\pm}}=(d_{Ax}\pm d_{Bx})/\sqrt{2}$ ($A,B$ the layer index, $x\equiv d_{x^{2}-y^{2}}$, $z\equiv d_{z^{2}}$). Compared to the high-pressure bulk crystal, the thin film has two main differences in the TB model parameters that significantly affect the bands and Fermi surface topology. First, the inter-layer intra-$z$-orbital hopping $t_{\perp,\text{0.5UC}}^{z}$=-0.439 eV is nearly 30% (0.2 eV) smaller in magnitude than $t_{\perp,\text{HP}}^{z}$=-0.635 eV. Second, the crystal-field splitting $\Delta E_{\text{0.5\text{UC}}}=\epsilon_{x}-\epsilon_{z}$ is enhanced to 0.519 eV, roughly $\sim$40 % larger than $\Delta E_{\text{HP}}=0.367$ eV. Smaller $t_{\perp}^{z}$ means smaller bonding-anti-bonding splitting, which leads to the coexistence of $\delta$ electron pocket (located at $\Gamma$ point) formed by the $z_{-}$ band and the $\gamma$-hole pocket (located at $M$ point) formed by $z_{+}$ band, as shown in Fig. 2c-d for both hole-doped (with 0.2 hole doping per Ni) and un-doped scenarios. This coexistence, however, does not happen in both high-pressure[11] and ambient pressure bulk crystal[36].
We now compare the Fermi surfaces obtained from DFT+$U$ and ARPES experiments at both undoped ($n=1.5$) and hole-doped ($n=1.3$, as estimated in ARPES experiments[37]) scenario. Significant discrepancies between DFT+$U$ results and ARPES data are apparent. For instance, the distance $d$ (marked by the red arrow in Fig. 2c) between two Fermi surface branches of the $\beta$ band near $(\pi,0)$ point is much smaller in AREPS than that in DFT results. The $\gamma$ pocket around $(\pi,\pi)$ seen in ARPES is much smaller than that in the DFT+$U$, and its diffuse nature may indicate that the $d_{z^{2}}$ states are much less coherent than the $d_{x^{2}-y^{2}}$ states. We will demonstrate that these discrepancies arise from the incomplete treatment of electron correlations in the DFT+U approach.
The second part is the rotation-invariant Kanamori-type multi-orbital on-site interaction. Realistic intra-orbital Coulomb interaction $U_{\mathrm{RPA}}\approx$ 3.77 eV and Hund’s coupling $J_{\mathrm{RPA}}\approx$0.56 eV (see Table 3) are obtained according to the constrained random phase approximation (cRPA)[38], which counts in the screening effect from the rest energy bands out of the low-energy $e_{g}$ subspace. As non-local interaction $V$ can screen $U$ further[17] while $J$ unscreened, we tune $U$ close to $U_{\mathrm{RPA}}$ with fixed $J=J_{\mathrm{cRPA}}$= 0.56 eV to mimic this screening effect and also systematically investigate the effect of $U$ on the correlated electronic structure. By comparing the electronic structure calculated from CDMFT and that of the ARPES experiment, we find that the best fitting happens at $U\sim 3.6$ eV (will be shown below). In this regime, CDMFT shows clear quasi-particle (QP) bands but with some small QP weight ($Z_{d_{z}^{2}}\sim 0.2\ll 1$), which suggests that the system is away from a doped Mott insulator but still have a significant correlation effect. In other words, our CDMFT results indicate that the thin film is an intermediate correlated system.
CDMFT Studies
The main results of CDMFT are summarized in Fig.3, which in detail demonstrate how the physical quantities like quasi-particle weight, occupancy, effective energy level, spectral function, and Fermi surfaces vary as one increases $U$. The calculations are performed at $T$=200 K and $n$=1.3 per site, which is close to the filling 1.28 estimated in the ARPES experiment[37]. More results on doping and temperature dependence can be found in SM.3, especially the results at lower temperatures like $T=$ 50 K, which only have quantitative differences from the results at $T=$ 200 K.
As one increases $U$, we observe an increase in effective energy splitting $\epsilon^{\text{eff}}_{z_{-}}-\epsilon^{\text{eff}}_{z_{+}}$ between the $z_{+}$ and $z_{-}$ orbitals as $U$ increases, where $\epsilon^{\text{eff}}_{\alpha}=\epsilon_{\alpha}+\mathrm{Re}\Sigma{\alpha(i0^{+})}-\mu$. Thus the effective energy level $\epsilon^{\text{eff}}_{z_{-}}$ ($\epsilon^{\text{eff}}_{z_{+}}$) of the $z_{-}$ ($z_{+}$ ) orbital is pushed upward (downward) as seen in Fig. 3a. For $U\leq$ 3.0 eV, $\epsilon^{\text{eff}}_{z_{\pm}}$ deviates slowly from its bare value $\epsilon_{z_{\pm}}$ (dashed lines). The deviation increases rapidly when $U\gtrsim 3.6$ eV, as indicated by the arrows. This is consistent with the picture of strong inter-layer electronic correlations between the $z$ orbitals[25], which becomes more prominent as $U$ increases. On the contrary, $\epsilon^{\text{eff}}$ for the $x$ orbitals remains gradually increasing even for larger $U$. $U=3.6$ eV could be regarded as a characteristic interaction strength where the correlation effect becomes significant.
The $U$-dependent orbital occupancy per spin is given in Fig. 3c. There is a self-doping effect with the electrons transfer from the $x$ orbitals to the $z$ orbitals (note that $n_{d}$=1.3 fixed). As $U$ is tuned from 2 to 3.77 eV, the total filling in the $z$ orbitals increases from $\sim$0.76 to $\sim$0.9 while the total filling in the $x$ orbitals decreases from $\sim$0.54 to $\sim$ 0.4. The quasi-particle weight $Z$ of the $x_{\pm}$ orbitals ranges from $\sim$0.7 to $\sim$0.5 as $U$ is tuned down, which results from a small filling per spin ($\sim$ 0.3 to $\sim$0.2). The $Z$ of $z_{\pm}$ orbitals decreases from $\sim$0.45 to $\sim 0.2$ for the $z_{+}$ orbital and to $\sim$ 0.08 for $z_{-}$ orbital, showing that $z$ orbitals are much more correlated than the $x$ orbitals. The reason for this is that the self-doping effect from $x$ to $z$ orbital makes the filling in each $z$ orbital approach to half filling per spin (from $\sim$0.38 at $U$=2 to $\sim$0.45 at $U=3.77$.)
Associated with the enhanced effective level splitting as $U$ increases, the ${\bf k}$ spectral functions $A({\bf k},\omega)$ and Fermi surfaces evolve correspondingly. From panel 3e-h, one can see that both $\delta$-band at $\Gamma$ point and $\gamma$-band at $M$-point at low-energy become more flat and gradually fade away (being less coherent) from the Fermi energy. First, as mentioned before, there is a downward or upward shift of $\epsilon^{\text{eff}}_{z_{\pm}}$. Second, the quasi-particle spectral weight (see Fig.3b) decreases, which makes especially the $\gamma$ band more flat near the Fermi energy. Furthermore, the scattering rate is enhanced when $U$ increases (not shown). As a result, the Fermi surface of $\delta$ electron pocket gradually disappears, as seen from Fig.3i-l . Simultaneously, the $\gamma$ hole pocket also shrinks in size and becomes less coherent (as evidenced by its decreased intensity in colormap). At $U=3.77$, the $\gamma$ band becomes flat near $M$ point, and its band-top touches the Fermi energy, which gives rise to the Fermi surface shown in Fig. 3n. The $A({\bf k},\omega)$ and Fermi surfaces at $U=3.6$ to $U=3.77$ eV shown in Fig. 3 quantitatively match all the key features seen in the ARPES Fermi surfaces, as indicated by the white dashed lines in Fig. 3l. (Doping and temperature dependence of the physical quantities in CDMFT are presented in SM3.). The agreement with ARPES data originates from the enhanced effective level splitting for $U\gtrsim 3.6$ eV, suggesting that the thin film is in an intermediate coupling regime.
Pairing Mechanism
The superconducting pairing symmetry is studied with a modified RPA approach, where the quasiparticle Hamiltonian from the CDMFT is taken as an input (see Method and SM.4 for details). We assume that effective interaction used in RPA is in the same form as the one in Eq. 9 and we fix $J_{\textrm{eff}}=U_{\textrm{eff}}/6$ in the calculations. The effective interaction strength $U_{\textrm{eff}}$ is a tuning parameter. It could be regarded as the strength of some kinds of residual interactions renormalized by a factor $Z^{2}$, where $Z$ is the quasi-particle weight. Since $Z$ of $d_{z^{2}}$ orbital is quite small ($\sim 0.18$ for $U=3.6$ eV), the effective interaction strength $U_{\textrm{eff}}$ should also be very small.
The RPA results for $U=3.6$ eV and $n=1.3$ are summarized in Fig. 4. In Fig. 4a, we show the typical RPA-renormalized spin susceptibility. The dominant wavevector $\mathbf{Q}_{1}$ corresponds to the nesting between the $\beta$-pocket and $\gamma$-pocket as shown in Fig. 4c. In Fig. 4b, we show the dependence of effective pairing strength[39] $\lambda$ on $U_{\textrm{eff}}$ for various pairing symmetries. It is clear that larger $U_{\textrm{eff}}$ leads to stronger superconducting instability. The leading pairing symmetry is always a $s^{\pm}$-wave, where the gap function on the Fermi surface is depicted in Fig. 4c. The coincidence of $\mathbf{Q}_{1}$ in Fig. 4a and the one connects the largest gap in magnitude in c supports the spin-fluctuations-mediated pairing mechanism. In real space, the strongest pairing is between the inter-layer $d_{z^{2}}$ orbitals as shown in Fig. 4d, which is consistent with the fact that the Fermi surface patches connected by $\mathbf{Q}_{1}$ on $\beta$ and $\gamma$ bands shown in Fig. 4c are mainly formed by $d_{z^{2}}$ orbitals. We also perform similar calculations for $3\leq U\leq 3.77$, and $s^{\pm}$-wave pairing is always dominant (see SM.4 for more details). All in all, the most possible pairing symmetry is $s^{\pm}$-wave pairing for realistic parameters.
Discussion
First, we discuss the extent of correlation effects in the bilayer nickelate films. The on-site Coulomb repulsion from cRPA is $U\sim 3.77$ eV, consistent with the typical value of $U=3-4$ eV for the transition metals. Due to the strong interlayer coupling, the multi-orbital tight-binding model involves four bands with a total bandwidth around 4 eV, which is comparable to the on-site Hubbard $U$. The situation is similar to the one in the iron-based superconductor[40, 41, 42, 43], suggesting that the bilayer nickelate systems also lie in an intermediate correlation regime.
Although the correlation is moderate, its effects are non-negligible. The weak-coupling fluctuation-exchange (FLEX) approximation[44, 45, 29] at the same filling $n=1.3$ fails to reproduce ARPES Fermi surfaces. As a weak coupling approach, FLEX calculations are restricted to $U\leq$1.8 eV, beyond which the method may lose convergence. In the weak $U$ region, the FLEX can reproduce the Fermi surface obtained in CDMFT for $U\leq$ 3.0 (see SM.5). However, FLEX fails to see a Fermi surface topology similar to Fig. 3k obtained in CDMFT at $U=3.6$ eV. This indicates that FLEX can not describe the strong interlayer correlation which plays a crucial role in the electronic structure of the bilayer nickelate thin films.
Next, we discuss the comparison between strained thin films and pressurized bulks. The coupling between interlayer $d_{z^{2}}$ orbitals ($t_{\perp}^{z}/J_{\perp}^{z}$) are believed to be crucial in the pressurized bulks[13, 29, 19]. However, the thin films have a much longer interlayer distance than the that in the pressurized bulks, which leads to much weaker interlayer couplings, say about $30\%$ lower for $t^{z}_{\perp}$ and $50\%$ lower for $J^{z}_{\perp}$ (since it proportion to $(t^{z}_{\perp})^{2}$). It would be a challenge for the theories strongly depending on $t^{z}_{\perp}$ and/or $J^{z}_{\perp}$ to explain the same order-of-magnitude onset ${T_{\mathrm{C}}}$ of the two systems. In contrast, although DFT calculations suggest a distinct Fermi surface topology for the pressurized samples, the CDMFT-derived Fermi surface (which incorporates correlation effects and aligns with ARPES data) closely resembles that of the high-pressure case. Therefore, it would be much easier to understand why the thin films have same order-of-magnitude onset ${T_{\mathrm{C}}}$ with a much smaller $t^{z}_{\perp}$ given that the correlated effect has been included correctly.
At last, we would like to emphasize the quantitative effectiveness of the DFT+cRPA+CDMFT approach in the study of unconventional superconductors with weak to intermediate correlation effects. Simply with lattice parameter inputs from experiments, the ab initio calculations without parameter fine-tuning match ARPES results quantitatively. The following RPA or similar techniques may provide a qualitatively accurate description of the pairing symmetry on top of the precise correlated electronic structures. Therefore, this approach is feasible for studying superconductivity in weakly and moderately correlated systems within first-principles calculations. Its wide applicability in various materials may shift the paradigm of searching for unconventional superconductors.
{methods}
STEM experiment. All electron microscopy analyses were conducted using a FEI Themis Z transmission electron microscope operated at 200 kV. This microscope is equipped with a Cs Probe Corrector (DCOR) and a high-brightness field-emission gun (X-FEG) with a monochromator to enhance resolution and contrast. For scanning transmission electron microscopy (STEM) high-angle annular dark field (HAADF) and annular bright field (ABF) imaging, the inner and outer acquisition angles ($\beta$1 and $\beta$2) were set to 90 and 200 mrad for HAADF images, and 12 and 23 mrad for ABF images, respectively. The convergence angle was maintained at 25 mrad. The beam current was adjusted to approximately 40 pA for HAADF and ABF imaging. Cross-sectional STEM specimens were prepared using a FEI Helios G4 HX dual-beam focused ion beam/scanning electron microscope (FIB/SEM) system. The simulation of ABF images was performed using the multi-slice method implemented in the QSTEM software. The simulation parameters were carefully set to match those used during the electron microscopy characterization, ensuring that the simulated images accurately reflect the experimental conditions.